OpenStudy (anonymous):

The measure of one angle of a triangle is 40. The measure of the other two angles are in a ratio of 3:4

OpenStudy (anonymous):

180-40=140 140/7=20 3:4 *20 60:80 so the other 2 angles are 60 and 80

OpenStudy (anonymous):

total addition of angles = 180 so the remaining two angles= 180-40= 140 now let the common ratio =x so 3x+4x = 140 7x=140 x=20 so angles are 3x=3*20= 60 and 4x= 4*20 = 80